OpenStudy (anonymous):

what is the domain and range for 2x^2 - 8x +7

OpenStudy (anonymous):

Sorry, but before i answer, are you sure there's no "y" in it because with out a "y" there can be no range...

OpenStudy (monkey_sun):

lets assume you meant \[f(x)=2x ^{2}-8x+7\] The domain which is all values of x you are allowed to plug in to a function, is in this case all real #s or \[[-\infty,\infty]\] which means the set from negative infinity to positive infinity . now the range is the set of values you could possiably get out of that function when you plugged in those x's so if you try plugging in differant values for x you will soon see that none of the results will be less than -1. so the range is \[[-1,\infty]\] if you want to see this just plug the function into a graphing app or calculator and you will see a nice parabola with its bottom at y=-1 and goes up both ways forever. If you want to find that range without plugging in random values or graphing you will learn that in calculus.

OpenStudy (esshotwired):

dang you're on this monkey xD